OpenStudy (anonymous):

The manager of a gas station and convenience store tabulates the following facts: 40% of customers purchase gasoline (G). 36% of customers purchase cigarettes (C). 61% of customers purchase gasoline or a drink (D). 9% of customers purchase gasoline and cigarettes. 14% of customers purchase a gasoline and a drink. 15% of customers purchase cigarettes and a drink. 2% of customers purchase all three items. (a) If we randomly select a customer, what is the probability that she purchases gasoline or cigarettes? (b) If we randomly select a customer, what is the probability that

hartnn (hartnn):

P(A or B) = P(A)+ P(B)-P(A and B) here, A = gasoline, B =cigarette

hartnn (hartnn):

you know all 3, P(A),P(B),P(A and B) just plug in values!

OpenStudy (anonymous):

0.04+0.36-?

hartnn (hartnn):

40% = 0.4 and P(G and C) is given as % = 0.09

hartnn (hartnn):

9%

OpenStudy (anonymous):

what about P(A and B)

OpenStudy (anonymous):

oh I got it

hartnn (hartnn):

what you got as final answer ?

OpenStudy (anonymous):

what about b ?

hartnn (hartnn):

complete question ?

OpenStudy (anonymous):

i got 0.67

OpenStudy (anonymous):

is that right?

hartnn (hartnn):

lets check, P(G or C) = 0.4+0.36-0.09 = 0.76-0.09 = 0.67 yes, you are correct :)

OpenStudy (anonymous):

what about b?

hartnn (hartnn):

you haven't posted question for b

OpenStudy (anonymous):

If we randomly select a customer, what is the probability that she purchases a drink

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

?

hartnn (hartnn):

can you make a venn diagram for your question ?

OpenStudy (anonymous):

can you give me some idea?

OpenStudy (anonymous):



hartnn (hartnn):

yes^ from that can you find probability of drink ? just add all the values in D circle.

OpenStudy (anonymous):

Makes you wonder what the last 25% are doing. Using the washroom or casing the joint.

OpenStudy (anonymous):

0.35?

OpenStudy (anonymous):

That looks good to me.

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

t more quick questions

OpenStudy (anonymous):

Which of the following statements is true? G and D are independent. G and C are independent. C and D are independent. None of the three pairs of events are independent.

OpenStudy (anonymous):

You might want to go over how to do that diagram. That's a fundamental skill for these types of questions.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

(d) Which of the following statements is true? C and D are mutually exclusive. G and D are mutually exclusive. G and C are mutually exclusive. None of the three pairs of events are mutually exclusive.

hartnn (hartnn):

mutually exclusive events have intersection = 0 or if P(A and B)=0, then A, B are mutually exclusive.

hartnn (hartnn):

so, do you find any such pair here ?

OpenStudy (anonymous):

G and D are mutually exclusive.

OpenStudy (anonymous):

?

hartnn (hartnn):

G and D = 12+2 , its not 0

OpenStudy (anonymous):

None of the three pairs of events are mutually exclusive.

hartnn (hartnn):

yeah.

OpenStudy (anonymous):

what about the previous one about independent?

hartnn (hartnn):

for independent , P(A and B) = P(A)*P(B) so, you need to check this for all 3 pairs.

OpenStudy (anonymous):

pg and pd , i got 0.14

hartnn (hartnn):

how ?

OpenStudy (anonymous):

P(G or D)=0.61

OpenStudy (anonymous):

Pg is 0.4

OpenStudy (anonymous):

14% of customers purchase a gasoline and a drink P(g and d)so it;s going to be 0.14

hartnn (hartnn):

'or' ?? you need to use 'and'

hartnn (hartnn):

P(g and d)so it;s going to be 0.14 is correct

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

G and C we have 0.65

hartnn (hartnn):

is that = p(g)*p(d) ??

OpenStudy (anonymous):

oh no

hartnn (hartnn):

then G and D are not independent.

OpenStudy (anonymous):

0.144?

hartnn (hartnn):

similar check for other 2 pairs.

OpenStudy (anonymous):

why they are not independent?

OpenStudy (anonymous):

0.216

OpenStudy (anonymous):

so None of the three pairs of events are independent.

hartnn (hartnn):

yeah

OpenStudy (anonymous):

ok last question, i actually got the answer could you help me to check it?

hartnn (hartnn):

yes?

OpenStudy (anonymous):

What is P(C Dc), i.e., the probability that a randomly selected customer purchases cigarettes but not a drink?

OpenStudy (anonymous):

i got 0.21

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

thank you very much?

hartnn (hartnn):

welcome?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

for this part

OpenStudy (anonymous):

Which of the following statements is true? G and D are independent. G and C are independent. C and D are independent. None of the three pairs of events are independent.

OpenStudy (anonymous):

answer is not None of the three pairs of events are independent

OpenStudy (anonymous):

?

hartnn (hartnn):

i didn't calculate for all 3, but looking at numbers, i think its None...

OpenStudy (anonymous):

yeah, but its not.

OpenStudy (anonymous):

can you help me to check it please?

hartnn (hartnn):

i still get none :( lets have a 2nd opinion by @tcarroll010

OpenStudy (anonymous):

help me on this one

OpenStudy (anonymous):

Which of the following statements is true? G and D are independent. G and C are independent. C and D are independent. None of the three pairs of events are independent.

OpenStudy (anonymous):

im not sure, what are independent

OpenStudy (mertsj):

Independent means that the occurrence of one does not affect the probability of the other.

OpenStudy (mertsj):

Is there just one answer?

OpenStudy (anonymous):

in this one, what are independent

OpenStudy (anonymous):

yeah, they are multiple choice

OpenStudy (mertsj):

Then I guess I would choose G and C